# The cognitive triple-slit experiment

### v2

@@ 6 Concluding remarks
In our cognitive test, we observed some interesting similarities with physics experiments. For example, it was not obvious that additional fringes would form, in addition to those between the two slits. This means that there is enough richness in human cognitive processes to produce complex patterns that resemble those produced by interference phenomena in quantum physical processes. However, some important differences were also observed, which were nevertheless expected.
@@
Our design was a sort of hybrid between a purely cognitive experiment, such as the one conducted by Hampton [12] (described in Section 1) and a physics experiment. Our motivation in proceeding in this way was to test one of the explanations offered by the conceptuality interpretation, consisting in explaining the interference fringes, in particular the central one, as regions of actualization of a situation of doubt about the slit from which the quantum entity emerged.
Before commenting further on these similarities and differences, let us model Hampton’s data in a similar way to how we did in our experiments. To obtain the complex birefringence pattern shown in Figure 2, Aerts represented the concepts Fruits, Vegetables and Fruits or vegetables by unit vectors $|F\rangle$, $|V\rangle$ and $|FV\rangle={1\over\sqrt{2}}(|F\rangle+|V\rangle)$, respectively. He then chose them so that the wave functions $\langle x|F\rangle$ and $\langle x|F\rangle$ would be two-dimensional, with a Gaussian real part. Placing all the exemplars on a two-dimensional plane, such that for their coordinates the modules squared $|\langle x|F\rangle|^{2}$, $|\langle x|V\rangle|^{2}$ and $|\langle x|FV\rangle|^{2}$ properly modeled the experimental probabilities, he then obtained an interference-like pattern reminiscent of the phenomenon of birefringence [11]; see Figure 2.
The use of Gaussians requires at least two dimensions in order to adequately distribute the different datapoints. However, this does not mean that it is not possible to also describe Hampton’s data using a one-dimensional screen, as we did in our experiment. For this, one has to position the exemplars in such a way that two clear peaks appear for the two single-slit experiments relating to the concepts Fruits and Vegetables, centered on the exemplars that obtained the highest score. Obviously, this requires giving up perfectly symmetrical Gaussian distributions. Once this is done, one can deduce the two-slit curve for the combination Fruits or Vegetables, as shown in the left graph of Figure 10.
Figure 10: (color online) Left: A continuous approximation of the discrete single-slit functions $P_{F}(x)$ (purple color) and $P_{V}(x)$ (green color), and corresponding double-slit function $P_{FV}(x)$ (brown color), which is the one-dimensional equivalent of the two-dimensional pattern exhibited in Figure 2. Right: A continuous approximation of the discrete single-slit functions $P(x|1)$ (purple color) and $P(x|3)$ (green color), and corresponding double-slit function $P(x|13)$ (brown color).
@@
It is important to note that although the strong variability of $P_{FV}(x)$ can be interpreted as the effect of an interference phenomenon, there are no actual waves from which the pattern would originate. Quantum wave functions should not be interpreted as real waves either, being clear that they exist in a configuration space that is distinct from the physical space, whose number of dimensions grows proportionally to the number of entities present. According to the conceptuality interpretation, what we interpret in quantum mechanics as interference phenomena, associated with a wave-like nature of the entities under study, would more fundamentally be the result of a variability in the meanings associated with the different outcomes, when these are organized and represented according to a certain logic. In other words, by recognizing that there is a more fundamental level governed by meaning and cognitive processes, it would be possible to overcome the wave-particle duality and understand the formation of patterns whose origin would not be wave-like but genuinely cognitive-like.
@@
One possible objection to the conceptuality interpretation, in relation to its interpretation of the double-slit and triple-slit experiments, is that the patterns observed in physics laboratories are usually very regular, whereas those obtained by analyzing cognitive experiments à la Hampton present much more variability. It is not difficult to respond to this type of criticism. If we truly live in a pancognitivist reality, as the conceptuality interpretation suggests, then the structures of meaning that have emerged within our human culture represent a very recent episode and have not yet had the opportunity to stabilize and organize themselves in the same way as the much older cognitive structures that supposedly form our physical universe.
@@
In the case of human cognition, things are also complicated by the fact that there are at least two lines to go from abstract to concrete concepts, while for entities in the material world only one of these lines would be active (see [15, 18] for an in-depth discussion of this point). Cognitive experiments should therefore only be considered qualitatively similar to experiments in physics laboratories, without expecting them to correspond to the latter in every possible aspect. In any case, even between similar cognitive experiments, significant structural differences are to be expected, as is the case between ours and Hampton’s experiments, where for example in one there is a relatively intense central peak between the two slits, whereas in the other such a peak is absent; see Figure 10.
@@
In conclusion, our cognitive triple-slit experiment adds to our understanding of human cognitive processes in situations of uncertainty and provides further confirmation of the explanatory power contained in the conceptuality interpretation, which is still under investigation by our Brussels group. The observed patterns, structurally reminiscent of those found in quantum interference experiments, speak to a deeper interpretative layer, one where meaning and information play the role traditionally attributed to physical interactions. In this sense, our results not only support the idea that cognition can manifest interference-like effects, but also reinforce the plausibility of a pancognitivist reality, in which physical and human cognitive processes are governed by analogous principles operating at different levels of abstraction.
@@ Appendix A The interference fringes in the double and triple-slit experiment
In this appendix, we show how to identify (with the necessary simplifications) the positioning of the different interference fringes in the double-slit and triple-slit experiments, as well as the typical width of the peak in a single-slit situation, in the Fraunhofer (far-field) regime. Let us start with the situation with two slits, without considering the width of the slits, which will be initially disregarded. Let $L$ and $d$ be the distances between the detector screen and the slit plate, and between the two slits, respectively; see Figure 11.
Figure 11: Left: the distances $L_{1}$ and $L_{2}$ from the two slits to point $x$ on the detector screen, and the difference in length of the two paths $\Delta_{12}=L_{1}-L_{2}$. The drawing also shows the distance $L$ between the detector screen and the slit plate, and the distance $d$ between the two slits. Right: the distances $L_{1}$, $L_{2}$ and $L_{3}$ from the three slits to point $x$ on the detector screen, and the differences in length of the three paths $\Delta_{12}=L_{1}-L_{2}$ and $\Delta_{23}=L_{2}-L_{3}$.
@@
Clearly, $L_{1}=\sqrt{L^{2}+\left(x+{d\over 2}\right)^{2}}$ and $L_{2}=\sqrt{L^{2}+\left(x-{d\over 2}\right)^{2}}$, hence, assuming that $x,d\ll L$, we have the approximation $\Delta_{12}=L_{1}-L_{2}\approx{xd\over L}$. The constructive interferences are characterized by the waves arriving in phase, therefore by the condition $\Delta_{12}=n\lambda$, where $\lambda$ is the incoming wavelength and $n\in\mathbb{Z}$. Destructive interferences, on the other hand, are characterized by a phase opposition condition, therefore $\Delta_{12}=\left(n+{1\over 2}\right)\lambda$. Hence, the positions of the constructive and destructive interference peaks are given by $x=n{\lambda L\over d}$ and $x=\left(n+{1\over 2}\right){\lambda L\over d}$, respectively. Therefore, if we neglect diffraction phenomena due to the finite size of the slits, all fringes have the same intensity and are separated from each other by the same distance ${\lambda L\over d}$, which is directly proportional to the wavelength $\lambda$. However, it is important to note that, in the construction of a wave packet, which involves the superposition of numerous wavelengths, only for the central fringe ($n=0$) will the constructive interference point be the same for all wavelengths. Therefore, the central fringe will necessarily be the brightest.
Let us now consider the situation with three slits; see Figure 11. Reasoning in a similar way as before, and assuming that also the approximation $d\ll x$ holds (this means that our reasoning will no longer be valid for the central fringes), we have $\Delta_{12}=L_{1}-L_{2}\approx{xd\over L}$, $\Delta_{23}=L_{2}-L_{3}\approx{xd\over L}$ and $\Delta_{13}=L_{1}-L_{3}=L_{1}-L_{2}+L_{2}-L_{3}=\Delta_{12}+\Delta_{23}\approx{2xd\over L}$.
@@
So, if we call $\psi_{1}$, $\psi_{2}$ and $\psi_{3}$ the three amplitudes at point $x$, associated with the three slits, the intensity $I(x)$ of their superposition is given by the modulus squared of their sum: $I(x)=|\psi_{1}+\psi_{2}+\psi_{3}|^{2}=I_{0}|1+e^{ik\Delta_{12}}+e^{ik\Delta_{13}}|^{2}$, where $k={2\pi\over\lambda}$ is the angular wave number. Using the previous approximations, we can write: $I(x)\approx I_{0}|1+e^{ik{xd\over L}}+e^{i2k{xd\over L}}|^{2}$. Clearly, the peaks of maximum intensity are exactly in the same positions as the two-slit problem. However, the absolute minima are not. Indeed, they correspond to the values $x=\left(n\pm{1\over 3}\right){\lambda L\over d}$, for which one can easily check that $I(x)=0$. In addition to the absolute maximums and minimums, in the three-slit problem there are also local peaks (fringes of lesser intensity); see Figure 11. These appear exactly where, in the two-slit problem, the absolute minima were found. Indeed, for points $x=n{\lambda L\over d}$, we have $I(x)=I_{0}$. The intensity of these peaks is therefore one third of that of the maximum peaks.
Note that the approximation $d\ll x$ does not allow to determine the condition for the central fringe, or fringes close to it. However, for $x$ small we can use the approximations $\Delta_{12}\approx{d\over L}(x-{d\over 2})$ and $\Delta_{23}\approx{d\over L}(x+{d\over 2})$. Cleary, for symmetry reasons, $\Delta_{13}=0$, for all $\lambda$, but we also have $\Delta_{12}\approx 0$ and $\Delta_{23}\approx 0$, since $d\ll L$, which explains why the central fringe remains the most intense one also in the situation with three slits.
@@
[…]

### v3

@@ 6 Concluding remarks
In our cognitive experiment, we observed some interesting similarities with physics experiments. For example, it was not obvious that additional fringes would form, in addition to those between the two slits. This means that there is enough richness in human cognitive processes to produce complex patterns that resemble those produced by interference phenomena in quantum physical processes. However, some important differences were also observed, which were nevertheless expected.
@@
Our design was a sort of hybrid between a purely cognitive experiment, such as the one conducted by Hampton [12], described in Section 1, and a physics experiment. Our motivation in proceeding in this way was to test one of the explanations offered by the conceptuality interpretation, consisting in explaining the interference fringes, in particular the central one, as regions of actualization of a situation of doubt about the slit from which the quantum entity emerged.
@@
It is important to note that although the strong variability of a fringe-like pattern can be interpreted as the effect of an interference phenomenon, there are no actual waves from which such pattern would originate in a cognitive experiment. Quantum wave functions should not be interpreted as real waves either, being clear that they exist in a configuration space that is distinct from the physical space, whose number of dimensions grows proportionally to the number of entities present. According to the conceptuality interpretation, what we interpret in quantum mechanics as interference phenomena, associated with a wave-like nature of the entities under study, would more fundamentally be the result of a variability in the meanings associated with the different outcomes, when these are organized and represented according to a certain logic. In other words, by recognizing that there is a more fundamental level governed by meaning and cognitive processes, it would be possible to overcome the wave-particle duality and understand the formation of patterns whose origin would not be wave-like but genuinely cognitive-like.
@@
One possible objection to the conceptuality interpretation, in relation to its interpretation of the double-slit and triple-slit experiments, is that the patterns observed in physics laboratories are usually very regular, whereas those obtained by analyzing cognitive experiments à la Hampton, or like ours, present much more variability. It is not difficult to respond to this type of criticism. If we truly live in a pancognitivist reality, as the conceptuality interpretation suggests, then the structures of meaning that have emerged within our human culture represent a very recent episode and have not yet had the opportunity to stabilize and organize themselves in the same way as the much older cognitive structures that supposedly form our physical universe.
@@
In the case of human cognition, things are also complicated by the fact that there are at least two lines to go from abstract to concrete concepts, while for entities in the material world only one of these lines would be active (see [16, 19] for an in-depth discussion of this point). Cognitive experiments should therefore only be considered qualitatively similar to experiments in physics laboratories, without expecting them to correspond to the latter in every possible aspect. In any case, even between comparable cognitive experiments, significant structural differences are to be expected, as is the case between ours and Hampton’s experiments, where for example in one there is a relatively intense central peak between the two slits, whereas in the other such a peak is absent (see Figure 3).
@@
In conclusion, our cognitive triple-slit experiment adds to our understanding of human cognitive processes in situations of uncertainty and provides further confirmation of the explanatory power contained in the conceptuality interpretation. The observed patterns, structurally reminiscent of those found in quantum interference experiments, speak to a deeper interpretative layer, one where meaning and information play the role traditionally attributed to physical interactions. In this sense, our results not only support the idea that cognition can manifest interference-like effects, but also reinforce the plausibility of a pancognitivist reality, in which physical and human cognitive processes would be governed by analogous principles operating at different levels of abstraction.
@@ Appendix A The interference fringes in the double and triple-slit experiment
In this appendix we derive the positions of interference fringes in double- and triple-slit experiments, and the typical peak width in the single-slit case, within the Fraunhofer regime. We begin with the two-slit setup, initially neglecting slit width. Let $L$ and $d$ be the distances between the detector screen and the slit plate, and between the two slits, respectively; see Figure 11.
Figure 11: Path lengths from the slits to a screen point $x$, and their differences, for the two-slit (left) and three-slit (right) cases.
@@
Clearly, $L_{1}^{2}=L^{2}+\left(x+{d\over 2}\right)^{2}$ and $L_{2}^{2}=L^{2}+\left(x-{d\over 2}\right)^{2}$, hence, assuming $x,d\ll L$, we have $\Delta_{12}=L_{1}-L_{2}\approx{xd\over L}$. The constructive interferences are characterized by the waves arriving in phase, therefore by the condition $\Delta_{12}=n\lambda$, where $\lambda$ is the incoming wavelength and $n\in\mathbb{Z}$. Destructive interferences, on the other hand, are characterized by a phase opposition condition, therefore $\Delta_{12}=\left(n+{1\over 2}\right)\lambda$. Hence, the positions of the constructive and destructive interference peaks are given by $x=n{\lambda L\over d}$ and $x=\left(n+{1\over 2}\right){\lambda L\over d}$, respectively. Therefore, if we neglect diffraction phenomena due to the finite size of the slits, all fringes have the same intensity and are separated from each other by the same distance $\lambda L\over d$}, directly proportional to the wavelength $\lambda$. Note that, in the construction of a wave packet, which involves the superposition of numerous wavelengths, only for the central fringe ($n=0$) will the point of constructive interference be the same for all wavelengths. Therefore, the central fringe will necessarily be the brightest.
@@
Consider now the situation with three slits; see Figure 11. Reasoning in a similar way as before, and assuming that also the approximation $d\ll x$ holds (this means that our reasoning will no longer be valid for the central fringe), we have $\Delta_{12}=L_{1}-L_{2}\approx{xd\over L}$, $\Delta_{23}=L_{2}-L_{3}\approx{xd\over L}$ and $\Delta_{13}=L_{1}-L_{3}=L_{1}-L_{2}+L_{2}-L_{3}=\Delta_{12}+\Delta_{23}\approx{2xd\over L}$. So, if we call $\psi_{1}$, $\psi_{2}$ and $\psi_{3}$ the three amplitudes at point $x$, associated with the three slits, the intensity $I(x)$ of their superposition is given by the modulus squared of their sum: $I(x)=|\psi_{1}+\psi_{2}+\psi_{3}|^{2}=I_{0}|1+e^{ik\Delta_{12}}+e^{ik\Delta_{13}}|^{2}$, where $k={2\pi\over\lambda}$ is the angular wave number. Using the previous approximations, we can write: $I(x)\approx I_{0}|1+e^{ik{xd\over L}}+e^{i2k{xd\over L}}|^{2}$. Clearly, the peaks of maximum intensity are exactly in the same positions as the two-slit problem. However, the absolute minima are not. Indeed, they correspond to the values $x=\left(n\pm{1\over 3}\right){\lambda L\over d}$, for which one can easily check that $I(x)=0$. In addition to the absolute maximums and minimums, in the three-slit problem there are also local peaks (fringes of lesser intensity); see Figure 11. These appear exactly where, in the two-slit problem, the absolute minima were found. Indeed, for points $x=n{\lambda L\over d}$, we have $I(x)=I_{0}$. The intensity of these peaks is therefore one third of that of the maximum peaks. Note that the approximation $d\ll x$ does not allow to determine the condition for the central fringe, or fringes close to it. However, for $x$ small we can use the approximations $\Delta_{12}\approx{d\over L}(x-{d\over 2})$ and $\Delta_{23}\approx{d\over L}(x+{d\over 2})$. Cleary, for symmetry reasons, $\Delta_{13}=0$, for all $\lambda$, but we also have $\Delta_{12}\approx 0$ and $\Delta_{23}\approx 0$, since $d\ll L$, which explains why the central fringe remains the most intense one also in the situation with three slits.
@@
$I(x)=|\int_{-{a\over 2}}^{a\over 2}\psi_{y}(x)dy|^{2}\approx I_{0}|\int_{-{a\over 2}}^{a\over 2}e^{i{kx\over L}y}dy|^{2}=I_{0}\left({\sin{kxa\over 2L}\over{kxa\over 2L}}\right)^{2}$ (9)